### v1

# Complex Magnetic Behavior in RuO2 Thin Films: Strain, Surface Effects, and Altermagnetic Signatures

Abstract
RuO2 has been proposed as a prototypical metallic $d$-wave altermagnet, a Néel-ordered compensated antiferromagnetic state exhibiting nonrelativistic momentum-dependent spin splitting; yet, its magnetic ground state remains controversial both theoretically and experimentally. Using comprehensive first-principles calculations, we investigate RuO2 thin films with (110), (100), and (001) orientations, both freestanding and supported on a TiO2 substrate. We show that magnetism in RuO2 thin films is highly fragile, strongly influenced by strain, surface orientation, and atomic relaxation, while also being highly sensitive to the choice of the Brillouin-zone integration scheme. Although tensile strain induces finite magnetic moments, none of the studied systems stabilizes a compensated antiferromagnetic state; hence, an altermagnetic ground state cannot be achieved. Instead, the magnetic response is characterized by pronounced layer- and site-dependent variations and incomplete moment compensation, resembling a ferrimagnetic-like state. Our results reconcile conflicting theoretical and experimental reports and underscore the sensitivity of RuO2 magnetism to structural and methodological details.
I Introduction
Finding and reliably identifying the true magnetic ground state of quantum materials remains one of the central challenges in condensed-matter physics and materials science [52, 4]. This task is particularly demanding in itinerant transition-metal oxides, where competing exchange interactions, electron itinerancy, and structural distortions often place the system close to multiple nearly degenerate magnetic instabilities [17, 49]. Even subtle perturbations, such as epitaxial strain, reduced dimensionality, surface termination, or details of electronic-structure methodology, can tip the balance between nonmagnetic, ferro-, ferri-, and antiferromagnetic states. Establishing the genuine magnetic ground state is, therefore, not only a fundamental problem but also a prerequisite for designing functional materials for spin-based technologies, where transport, symmetry, and magnetic order are intimately intertwined.
Recently, antiferromagnetic (AFM) materials exhibiting nonrelativistic, momentum-dependent spin splitting have been identified as a symmetry-distinct class of antiferromagnets beyond conventional spin-degenerate AFM systems [36, 1, 35, 57, 44, 47, 45, 29, 3]. In these systems, the spin splitting originates not from spin–orbit coupling but from the spin space/point group symmetry, which allows for exchange-driven lifting of spin degeneracy while preserving zero net magnetization. The resulting momentum-dependent spin splitting is constrained by crystal symmetry and can be classified according to its angular-momentum character in the electronic band structure. In particular, unconventional collinear AFM materials with even-parity spin splitting of $d$-, $g$-, or $i$-wave symmetry, collectively termed altermagnets, retain fully compensated magnetic moments in real space while exhibiting sign-changing spin polarization across the Brillouin zone [47, 45]. This symmetry-enforced structure of the spin splitting distinguishes it from conventional $s$-wave AFMs, in which combined inversion or parity and time-reversal symmetries protect spin degeneracy throughout momentum space.
Among these unconventional collinear AFMs, the $d$-wave class is particularly attractive because its symmetry allows for directional spin-polarized transport via the spin-splitter effect and spin splitting torque [16, 21]. Despite rapid progress in identifying candidate materials, many proposed altermagnets are insulating, which limits their immediate applicability in spin-transport devices. For practical applications, metallic or semiconducting systems are considerably more desirable, as they enable direct electrical control and detection of spin-polarized currents. RuO2 has been proposed as a prominent metallic candidate, with symmetry analysis predicting $d$-wave altermagnetic order. Early ab initio calculations [16, 46, 10] and some indirect quantum transport experimental observations [50, 10, 11, 18, 28, 60, 24, 23, 22] support this possibility. However, more recent high-resolution measurements, such as X-ray linear dichroism, nuclear magnetic resonance, spin-ARPES, and neutron diffraction; along with advanced first-principles studies, have challenged this picture, indicating that bulk RuO2 may, in fact, be nonmagnetic or, at most, extremely close to a magnetic [48, 40, 37, 20, 54]. This apparent contradiction has fueled an ongoing debate regarding the true magnetic ground state of RuO2 in bulk and thin films [8, 25, 33, 51, 27, 19, 55, 38, 31, 2, 7, 32, 30].
Ab initio methods based on density functional theory (DFT) are powerful tools for determining magnetic ground states. Standard DFT calculations without an on-site Hubbard $U$ correction predict a nonmagnetic ground state for bulk RuO2, whereas DFT+$U$ calculations with $U$ values of $U>0.6$ eV stabilize an altermagnetic state [27, 42, 48]. Although the applicability of DFT+$U$, which effectively enhances the localization of Ru 4dd electrons, may be questioned for a system with predominantly itinerant character, the fact that magnetism emerges already for relatively small $U$ values indicates that RuO2 lies in close proximity to a magnetic instability. This pronounced sensitivity implies that its electronic and magnetic properties can be strongly influenced by external perturbations, including strain, symmetry breaking at surfaces and interfaces, and chemical doping [42, 12, 56, 20, 53].
Recent experimental measurements have focused on the epitaxial growth of RuO2 thin films on TiO2 substrates, highlighting the importance of reduced dimensionality and substrate-induced effects on their magnetic properties. These developments underscore the need for a clear microscopic understanding of magnetism in RuO2 from first principles. Although several recent ab initio investigations have focused on the magnetism of RuO2 thin films [20, 6], significant gaps remain in our understanding of the detailed magnetic behavior of these systems.
In this study, we perform comprehensive DFT calculations to achieve a microscopic understanding of the magnetic behavior of RuO2 thin films with (110), (001), and (100) orientations, both under strain and in the strain-free limit. Furthermore, we investigate RuO2/TiO2 thin-film heterostructures with the same crystallographic orientations in order to elucidate the role of epitaxial coupling and interface effects.
Since the (110) orientation is the most thermodynamically stable surface for both TiO2 and RuO2 [9, 58], our analysis places particular emphasis on this configuration. For this orientation, we systematically assess the influence of Brillouin-zone sampling, structural relaxation, induced magnetization, and competing magnetic ordering patterns in order to obtain a robust and quantitatively reliable determination of the magnetic ground state.
Our calculations demonstrate that RuO2 single crystals and thin films remain non-magnetic in the absence of external perturbations. Magnetic order may develop only within a few layers near the interface under tensile strain. Nevertheless, even under these conditions, the system does not realize a compensated altermagnetic phase, but instead adopts a ferrimagnetic-like state with a finite net magnetization.
II Method and Structures
II.1 Ab initio details
We performed nonrelativistic, extitab initio density functional theory (DFT) calculations using the Quantum ESPRESSO package [14, 13, 15]. The electron–ion interactions were described using the Standard Solid-State Pseudopotentials (SSSP) library, PBEsol Efficiency version 1.3.0 [41]. For the exchange–correlation functional, we employed the PBEsol approximation [39], a revised form of the Perdew–Burke–Ernzerhof (PBE) generalized gradient approximation (GGA) that provides improved accuracy for the structural properties of solids and surfaces. A vacuum spacing of approximately 15 Å was introduced in all thin-film calculations to avoid spurious interactions between periodic images. Brillouin-zone integrations were performed using a uniform $k$-point grid corresponding to a reciprocal-space resolution of 0.10 Å-1. The thin-film structures were constructed using the CIF2Cell code [5], and the initial Quantum ESPRESSO input files were generated with the QE input generator [41]. The plane-wave kinetic energy cutoff was set to 40 Ry for the wave functions and 400 Ry for the charge density.
We investigate three RuO2 thin-film surface orientations: (110), (100), and (001), considered both as freestanding slabs and supported on TiO2 substrates (see Fig. 1). The RuO2/TiO2 heterostructures are constructed using the experimental crystal structure of TiO2. For the RuO2(110)/TiO2(110) system, recently studied experimentally, we examine RuO2 film thicknesses ranging from one to seven Ru atomic layers. For the freestanding RuO2(110) thin film, a 17-layer Ru slab is employed to ensure convergence with respect to slab thickness and to minimize spurious interactions between the two surfaces. For the supported RuO2(100)/TiO2(100) and RuO2(001)/TiO2(001) heterostructures, the RuO2 film thickness is fixed at 12 Ru atomic layers. In the case of freestanding RuO2(100) and RuO2(001) thin films, thicker slabs consisting of 33 and 35 Ru atomic layers, respectively, are used to achieve well-converged surface properties.
The RuO2/TiO2 heterostructure slabs were constructed by placing RuO2 layers symmetrically on both sides of the TiO2 substrate. This symmetric geometry avoids artificial effects associated with asymmetric slab models, where depositing RuO2 on only one side can introduce a net dipole moment across the slab. In such cases, spin-polarized calculations may yield spurious magnetic moments on the Ti atoms at the exposed surface. By eliminating the net dipole moment, the symmetric slab geometry suppresses these artifacts and ensures a physically meaningful description of the magnetic properties.
For the RuO2(110) thin film, three distinct $(1\times 1)$ terminations of the rutile structure have been reported in the literature [43]. In this work, we consider two of these terminations. The relative stability of these terminations depends on the oxygen chemical potential, reflecting the oxygen availability during growth or operation [43]. Under oxygen-rich conditions, the stoichiometric RuO2(110)–Obridge termination is the most stable and is widely regarded as the preferred configuration of the rutile (110) surface. Under oxygen-poor conditions, the surface instead favors the RuO2(110)–Ocus termination, characterized by coordinatively unsaturated Ru sites (cus). The third possible termination, RuO2(110)–Ru, is not considered here, as it is unstable even under strongly oxygen-deficient conditions. The two terminations examined in this work are illustrated in Fig. 1.
Since our explicit comparison of oxygen-rich and oxygen-poor terminations for the (110) orientation reveals no significant changes in the magnetic ordering, we confine our analysis of the (100) and (001) thin films to terminations without oxygen atoms atop the Ru sites.
The RuO2(110) orientation displays a repeating three-layer stacking sequence consisting of two oxygen-only layers and one mixed Ru–O layer. A similar three-layer periodicity is found for the RuO2(100) orientation, composed of two oxygen-only layers and one ruthenium-only layer. In contrast, for the (001) orientation, each atomic layer contains both Ru and O atoms. To adopt a unified notation for slab thickness, we define the thickness solely by the number of Ru-containing layers. For example, for the (110) surface, with its sequence of two O layers followed by one mixed Ru–O layer, we use the notation 1-L, 2-L, …, $n$-L instead of the conventional tri-layer notation (1-TL, 2-TL, …, $n$-TL).
Figure 1: Different surface orientations of RuO2 thin films considered in this work. For the RuO2(110)/TiO2(110) heterostructure, several film thicknesses were examined; the 5-L configuration is shown as a representative example. Blue and red spheres denote Ru atoms carrying different magnetic moments in the assumed AFM configurations discussed in the main text (labelled Ru1 and Ru2). Ti and O atoms are shown in green and white, respectively.
III Results
III.1 Magnetism and k-point sampling sensitivity
A recent study [42] examined the influence of strain on the magnetic properties of bulk RuO2, demonstrating that lattice expansion stabilizes magnetic order—an effect that has also been reported in other materials [59]. Furthermore, Ref. [42] showed that the calculated magnetic moment in bulk RuO2 is highly sensitive to the choice of k-point sampling, exhibiting pronounced fluctuations that were attributed to the fragile and unconventional nature of magnetism in this compound.
Here, we show that a significant part of these fluctuations originates from the specific Brillouin-zone integration scheme employed. To substantiate this point, we compute the absolute magnetization of RuO2 using the lattice parameters and atomic positions of rutile TiO2. This choice effectively applies a tensile strain to RuO2, under which the system stabilizes in an AFM ground state.
We tested several Brillouin-zone integration schemes across different k-point meshes, including smearing-based approaches such as Marzari–Vanderbilt–DeVita–Payne cold smearing [34] and Fermi–Dirac smearing. Because smearing methods depend on the choice of the broadening parameter ($\sigma$), we also employed the optimized tetrahedron method [26], which enables accurate Brillouin-zone integration without smearing. The resulting absolute magnetizations were then systematically compared across all integration schemes and k-point grids.
As shown in Fig. 2, the calculated magnetization is highly sensitive to both the choice of the smearing scheme and the value of the broadening parameter $\sigma$. For example, Fermi–Dirac smearing with $\sigma=0.01$$0.03$ Ry yields a nearly vanishing magnetization, whereas cold smearing under the same conditions produces a large absolute magnetization of approximately 1.0 $\mu_{\mathrm{B}}$. These results underscore the importance of carefully selecting the Brillouin-zone integration scheme and broadening parameter when investigating magnetism in RuO2.
In subsequent calculations, we performed structural relaxations using a small electronic broadening ($\sigma=0.00125$ Ry), which yields results nearly identical to those obtained with the optimized tetrahedron method. Final total-energy and magnetic-property calculations were then carried out using the optimized tetrahedron method to ensure an accurate treatment of Brillouin-zone integration.
Figure 2: Sensitivity of the Ru magnetic moment to k-point sampling in bulk rutile RuO2 under tensile strain. The calculations are performed for bulk RuO2 constructed using the lattice parameters and Wyckoff positions of rutile TiO2, which leads to a tensile strain and stabilizes an AFM state. k-point densities range from 0.05 Å-1 (corresponding to a $27\times 27\times 42$ Monkhorst–Pack mesh) to 0.30 Å-1 (corresponding to a $5\times 5\times 8$ mesh). (a) Ru magnetic moment as a function of the Monkhorst–Pack k-point grid calculated using Marzari–Vanderbilt–DeVita–Payne cold smearing for different smearing widths $\sigma$ (in Rydberg). Results obtained with the optimized tetrahedron method, which does not require a smearing parameter, are shown for comparison. (b) Same analysis performed using Fermi–Dirac smearing.
III.2 RuO2(110) thin films on TiO2(110)
Recent experimental investigations of the electronic band structure and transport properties of RuO2 have focused on RuO2(110) thin films grown on TiO2(110), owing to their favorable epitaxial matching and structural stability. We first investigate RuO2(110) thin films with thicknesses ranging from one to seven Ru-containing layers grown on a TiO2(110) substrate. Site-resolved magnetic moments are obtained by Bader partitioning of the spin density. In the AFM phase of bulk RuO2, the rutile structure contains two inequivalent Ru sites that form distinct magnetic sublattices, which we label Ru1 and Ru2. These sites are indicated in Fig. 1 by blue (Ru1) and red (Ru2) spheres.
Figure 3 presents the magnetic moments of Ru1 and Ru2 as a function of film thickness. In contrast to compensated AFM materials, the moments on Ru1 and Ru2 do not completely cancel. As a result, the RuO2(110) thin film grown on a TiO2(110) substrate develops a finite net magnetization, indicative of ferrimagnetic-like behavior.
Moreover, the magnetic moments exhibit a non-monotonic dependence on film thickness: the moments on both Ru1 and Ru2 oscillate as the number of Ru-containing layers increases, although without a strictly periodic pattern. These variations are most pronounced at the outermost surface layer and at the RuO2/TiO2(110) interface, underscoring the strong influence of surface and interface effects on the magnetic properties.
To further characterize the magnetic behavior of RuO2(110)/TiO2(110), we examine three key factors: the influence of the initial magnetic configuration, the role of atomic relaxation, and the impact of induced magnetic moments.
Because standard DFT calculations cannot systematically sample the full magnetic energy landscape, the choice of the initial magnetic configuration is particularly important for RuO2, where magnetic moments, if any, are intrinsically small. Exhaustive exploration of all possible magnetic initializations for different thicknesses is computationally prohibitive; we therefore perform a detailed analysis for a representative slab with a thickness of 3-L.
For this system, each layer contains two inequivalent Ru atoms (Ru1 and Ru2), yielding a total of six magnetic sites and ten distinct AFM configurations. We systematically examined all these AFM arrangements. Each configuration was initially imposed using magnetic constraints to stabilize the desired spin pattern; the constraints were subsequently released to allow full self-consistent relaxation. Among all tested configurations, the one initialized with antiparallel magnetic moments between Ru1 and Ru2 within each layer consistently relaxes to the lowest-energy state, characterized by opposite signs of the magnetic moments on Ru1 and Ru2, with incomplete compensation. The remaining AFM initializations either converge to higher-energy solutions or evolve into this same configuration upon relaxation. Moreover, for several slab thicknesses, calculations starting from a ferromagnetic state also converge to this energetically favorable AFM ordering. These results indicate that the ground state generally favors antiparallel alignment between Ru1 and Ru2 within each layer; accordingly, this antiparallel arrangement is adopted as the initial magnetic configuration in all subsequent calculations.
The only deviation from this trend occurs in the 6-L slab, where the Ru atoms in the third layer develop magnetic moments of the same sign on the two sublattices, despite being initialized in the constrained AFM configuration. Nevertheless, the magnetization in this layer remains strongly uncompensated, since the moment on Ru2 is significantly smaller than that on Ru1, resulting in a ferromagnetic alignment of the two sublattices, albeit with strongly unequal magnetic moments. The magnetic moments in the third layer are 0.33 $\mu\mathrm{B}$ on Ru1 and 0.04 $\mu\mathrm{B}$ on Ru2 for the bridge configuration, whereas for the cus configuration the corresponding values are 0.30 $\mu\mathrm{B}$ and 0.007 $\mu\mathrm{B}$, respectively.
Although full structural relaxations were carried out for all slab geometries, Fig. 4 shows that atomic relaxation has only a minor effect on the magnetic moments of the deeper RuO2(110) layers. Moreover, when the magnetic moment of the surface layer is artificially enhanced using a Hubbard $U$ correction, the resulting increase remains highly localized: the magnetic moments in the subsurface and deeper layers change only negligibly. This behavior demonstrates that the magnetization generated at the surface due to symmetry breaking [20] does not propagate into the deeper layers. Instead, the magnetic moments of the Ru atoms are primarily determined by strain-induced changes in their local atomic environment arising from interfacial tensile strain, rather than by long-range induction from the surface.
Figure 3: Magnetic moments of Ru1 and Ru2 as a function of layer index and film thickness. Filled and open symbols denote Ru1 and Ru2, respectively, for RuO2(110) thin films on a TiO2(110) substrate, with thicknesses ranging from 1-L to 7-L. For each thickness, magnetic moments are shown for all Ru layers within the slab. Data corresponding to the same film thickness are connected by lines of the same color to guide the eye.
Figure 4: Effect of atomic relaxation and induced magnetization on Ru magnetic moments in a 5-L RuO2(110) slab on TiO2(110). Magnetic moments of Ru1 (filled symbols) and Ru2 (open symbols) are shown as a function of the Ru layer index.
It is also instructive to compare the electronic structure of a compensated AFM RuO2(110)/TiO2(110) thin film with that of its true magnetic ground state. To this end, we calculate and compare the band structures of a 5-L RuO2(110) slab on a TiO2(110) substrate for two cases: (i) a constrained, fully compensated AFM configuration, which hosts nonrelativistic altermagnetic spin-split bands, and (ii) the unconstrained AFM ground state. In both cases, the non-relaxed geometry is employed to avoid additional effects arising from structural relaxation.
To make the comparison between the two configurations meaningful, we adjusted the fixed magnetic moments of the Ru atoms in the constrained AFM case so that their total absolute magnetization matches that of the unconstrained AFM ground state, where the magnetic moments vary from layer to layer (Fig. 3). Here, the total absolute magnetization is defined as $M_{\mathrm{abs}}=\int_{\mathrm{cell}}\left|n_{\mathrm{up}}(\mathbf{r})-n_{\mathrm{down}}(\mathbf{r})\right|\,d^{3}r$, where $n_{\mathrm{up}}(\mathbf{r})$ and $n_{\mathrm{down}}(\mathbf{r})$ are the spin-up and spin-down charge densities. By testing different values, we found that fixing the Ru moments to approximately $0.3~\mu_{\mathrm{B}}$ reproduces the total absolute magnetization of the unconstrained ground state. This value is therefore used in the constrained AFM calculation.
As shown in Fig. 5, the band structure of the magnetic ground state exhibits a substantially larger nonrelativistic spin splitting than that of the compensated AFM configuration. This enhancement arises because, in the true magnetic ground state, the Ru magnetic moments are not fully compensated, in contrast to the compensated AFM limit, where the net magnetization vanishes.
Figure 5: Nonrelativistic band structures of a 5-L RuO2(110) slab on a TiO2(110) substrate. The left panel shows the band structure of the unconstrained magnetic ground state, while the right panel shows the band structure obtained with Ru magnetic moments constrained to enforce a fully compensated AFM configuration. In both cases, the non-relaxed atomic structure is used. Red and blue lines denote spin-up and spin-down bands, respectively.
III.3 RuO2(100) and (001) thin films on TiO2(100) and (001)
We next consider the remaining low-index RuO2 surfaces, namely the (100) and (001) orientations. Owing to computational constraints, only one film thickness is examined for each case. For both orientations, we study a 12-layer RuO2 film grown on TiO2. Figure 6 shows the layer-resolved magnetic moments of the two inequivalent Ru sites (Ru1 and Ru2), before and after structural relaxation.
For both surfaces, atomic relaxation has a markedly stronger effect than in the (110) case. In the relaxed (001) and (100) films, magnetic moments in the outermost three layers are strongly suppressed, while deeper layers near the RuO2/TiO2 interface develop substantially larger local moments. By contrast, the unrelaxed films exhibit an almost ideal AFM arrangement, with Ru1 and Ru2 moments of similar magnitude and opposite sign; this ordering is significantly weakened upon relaxation.
Despite these differences, all systems, relaxed and unrelaxed, (001), (100), and (110), exhibit a characteristic layer-by-layer oscillation in the magnitude of the local magnetic moments. This behavior can be attributed to the metallic nature of the films: electronic states are confined between the surface–vacuum boundary and the buried interface, forming quasi-two-dimensional states analogous to a particle-in-a-box with finite potential barriers. The resulting quantum interference gives rise to oscillatory magnetic responses across the film thickness.
III.4 Intrinsic magnetism of freestanding RuO2 thin films
To further explore the role of the substrate, we also study freestanding RuO2 thin films with different surface orientations; see Fig. 1. Two classes of films are considered: (i) films constrained to the experimental lattice parameters of rutile TiO2, and (ii) films constrained to the experimental bulk lattice parameters of RuO2. In both cases, only the internal atomic coordinates are relaxed, while the lattice parameters are fixed. Since the in-plane rutile lattice constant of RuO2 is approximately 0.1 Å larger than that of TiO2, constraining the film to the TiO2 lattice imposes in-plane tensile strain. Therefore, adopting the TiO2 lattice parameters provides an effective way to apply biaxial tensile strain to RuO2 thin films.
Figure 7 shows the layer-resolved Ru magnetic moments for the (110), (100), and (001) orientations, both with and without tensile strain. In the unstrained (110) and (100) films, finite magnetic moments are largely confined to the first two surface layers and decay rapidly toward the bulk. The unstrained (110) result is consistent with recent reports (Ref. [20]).
In contrast, the (001) surface converges to a nonmagnetic solution, with no Ru atom exhibiting a discernible magnetic moment. This finding differs from recent reports of finite magnetization at the (001) termination [6]. To evaluate the robustness of our results, we carried out additional calculations using VASP with several different PBE pseudopotentials for Ru, namely Ru, Ru_pv, and Ru_sv, corresponding to the valence configurations $4d^{7}5s^{1}$, $4p^{6}4d^{7}5s^{1}$, and $4s^{2}4p^{6}4d^{7}5s^{1}$, respectively, as specified on the VASP website. All three yield the same outcome; for the unstrained (001) surface, the magnetic moments remain negligible ($10^{-6},\mu_{\mathrm{B}}$ per Ru atom). We therefore attribute the discrepancy to methodological differences, most likely related to the Brillouin-zone integration scheme, as our analysis indicates that the emergence of magnetism in RuO2 is highly sensitive to the details of $k$-space sampling.
Figure 6: Layer-resolved Ru magnetic moments for a 12-layer RuO2 slab with (100) and (001) orientations on a TiO2 substrate. Filled and open symbols denote Ru1 and Ru2 sites, respectively. Results are shown for both relaxed and unrelaxed atomic structures.
Figure 7: Layer-resolved Ru magnetic moments for freestanding RuO2 slabs with (110), (100), and (001) orientations. For each orientation, two unrelaxed reference structures are shown: one constrained to the experimental TiO2 lattice parameters (tensile strain) and one using the experimental bulk RuO2 lattice. Filled and open symbols denote Ru1 and Ru2 sites, respectively.
[…]
IV summary and conclusion
In this work, we have performed a comprehensive first-principles investigation of the magnetic properties of RuO2 thin films with (110), (100), and (001) orientations, both as freestanding slabs and in RuO2/TiO2 heterostructures. Motivated by the ongoing debate on the existence of $d$-wave altermagnetism in metallic RuO2, we systematically examined the role of epitaxial strain, surface termination, film thickness, structural relaxation, and Brillouin-zone integration schemes in determining the magnetic ground state.
Our calculations demonstrate that magnetism in RuO2 thin films is intrinsically fragile and highly sensitive to both structural and methodological details. In particular, we show that the calculated magnetic moments depend strongly on the choice of Brillouin-zone integration scheme and smearing parameters. When treated with sufficiently accurate integration, bulk RuO2 at its experimental lattice parameters remains nonmagnetic, consistent with several recent experimental reports that question robust long-range antiferromagnetism in this compound. This sensitivity places RuO2 in close proximity to a magnetic instability, where small perturbations can induce local moments.
Epitaxial tensile strain, mimicked by constraining RuO2 to the rutile TiO2 lattice parameters, induces finite Ru magnetic moments for all investigated surface orientations. However, in no case do we find stabilization of a fully compensated Néel-ordered AFM state and hence any altermagnetic behavior. Instead, the magnetic response is strongly layer- and site-dependent, with incomplete compensation between the two Ru sublattices (Ru1 and Ru2). In RuO2(110)/TiO2(110) heterostructures, the magnetization is dominated by surface and interface layers and exhibits pronounced oscillations with film thickness, consistent with quantum confinement and modified Ru–O hybridization at the boundaries. Even when initialized in various magnetic configurations, the system relaxes to states characterized by residual net magnetization, i.e., ferrimagnetic-like order rather than a symmetry-protected compensated AFM state.
For freestanding films, unstrained (110) and (100) slabs develop only weak surface-localized moments, while the unstrained (001) surface remains essentially nonmagnetic within numerical accuracy. Under tensile strain, all orientations develop enhanced magnetization, but the compensation between Ru1 and Ru2 moments remains incomplete. Notably, the (100) orientation exhibits the largest net magnetic moment under strain, yet still fails to realize a compensated altermagnetic phase. The (001) orientation shows the weakest magnetic response overall, even under tensile strain, making it the least favorable candidate for altermagnetic behavior among the studied surfaces.
By explicitly comparing the band structures of constrained compensated AFM configurations with those of the true magnetic ground states, we show that the larger nonrelativistic spin splitting observed in thin films originates from uncompensated magnetization rather than from a symmetry-enforced altermagnetic state. Therefore, any experimentally observed spin splitting or anomalous transport in RuO2 thin films must be carefully interpreted in light of possible ferrimagnetic-like surface or interface contributions, rather than being directly attributed to ideal $d$-wave altermagnetism.
Overall, our results reconcile seemingly conflicting theoretical and experimental findings by demonstrating that RuO2 resides at the brink of magnetism: small strain, reduced dimensionality, or methodological differences can induce local magnetic moments; yet none of the studied realistic thin-film geometries stabilize a fully compensated altermagnetic ground state. These findings underscore the importance of rigorous numerical convergence and detailed structural modeling when identifying candidate altermagnets, and they suggest that achieving robust altermagnetism in RuO2 may require additional symmetry engineering, chemical modification, or external tuning beyond simple epitaxial strain.
In addition, our findings call for a careful re-evaluation of recent anomalous Hall measurements in RuO2 thin films that have been interpreted within a fully compensated altermagnetic framework. Since our calculations reveal an unconventional ferrimagnetic-like order with finite net magnetization in strained and heterostructured geometries, conventional Berry-curvature-driven anomalous Hall contributions are expected. These epitaxial thin films grown on TiO2 can also host strong spin-orbit couplings. We therefore propose systematic thickness- and strain-dependent transport measurements to distinguish genuine symmetry-enforced altermagnetism from the unusual uncompensated magnetism predicted in this work.
ACKNOWLEDGMENTS
A.Q. has been supported by the Research Council of Norway through its Centers of Excellence funding scheme, Project No. 262633 “QuSpin” and FRIPRO with Project No. 353919 “QTransMag.”